three particles are located at the vertices of an equilateral triangle of side a .Each of the particle starts to move with a speed v,with the first particle heading continually for the second ,the second for the third and the third for the first .When would the particles meet each other?

Grade:12th pass

1 Answers

Arun
6 years ago
Let the particles are A, B, C

velocity of A is v along AB 

the velocity of B is along BC . its component along BA is vcos60 = v/2. thus the separation AB decreases at the rate

v+v/2=3v/2

since this rate is constt. the time taken in reducing the separation AB from d to 0 is  

t = d/(3v/2) =2d/3v

 

Regards

Arun (askIITians forum expert)